# Time-MMD: A New Multi-Domain Multimodal Dataset for Time Series Analysis

### v1

@@ 5 Broader Impacts
Beyond its efficacy in enhancing time-series forecasting accuracy through multimodality (Section 4), Time-MMD holds significant potential in advancing time series analysis across a wide spectrum. In the following section, we discuss how Time-MMD can transform conventional approaches, facilitate novel methodologies, and broadly impact the time series analysis domain.
@@
Multimodal Time-Series Imputation. Missing values in time series data, caused by sensor failures, system instability, or privacy concerns, pose a significant challenge in analysis. Conventional time-series imputation (TSI) methods (6; 14; 1; 2) often overlook valuable information captured in textual formats alongside the numerical data. For instance, incident reports, weather conditions, and special events can provide crucial context for imputing missing data points in traffic time series, but current methods fail to effectively incorporate this information. Time-MMD enables the integration of textual contextual information with numerical time series data, opening new avenues for multimodal time series analysis and enhancing imputation accuracy.
@@
Multimodal Time-Series Anomaly Detection. Detecting anomalies in time series data is crucial for identifying unusual patterns that may indicate faults, fraud, or other significant events (8; 3; 52). However, conventional anomaly detection methods (29; 5; 47; 34) are limited to expected pattern deviations in numerical data, overlooking valuable information in textual formats. For instance, news articles, social media posts, and market reports can provide critical context that influences financial market behavior and helps identify anomalies not evident from numerical data alone.
@@
Multimodal Foundation Time-Series Models The introduction of Time-MMD, a comprehensive text-numeric TS dataset, is expected to significantly advance multimodal-based TS methods, including the the development of multimodal foundation TS models. Time-MMD will facilitate further exploration with more complex and informative prompts, enhancing the performance and capabilities of fine-tuning methods. Additionally, it will spur research and development of multimodal models specifically tailored for TSA, an area with limited exploration compared to other domains like vision and video generation(35; 41; 46; 13).
@@ 7 Conclusion
Acknowledgements
This paper was supported in part by the NSF (Expeditions CCF-1918770, CAREER IIS-2028586, Medium IIS-1955883, Medium IIS-2106961, PIPP CCF-2200269, IIS-2008334, CAREER IIS-2144338), CDC MInD program, Meta faculty gifts, and funds/computing resources from Georgia Tech.
@@ Appendix B Limitations
Our work provides a comprehensive, high-quality multimodal time series dataset, but it still has limitations in terms of dataset diversity, as all the text data comes from English. We plan to extend Time-MMD to multilingual versions to better address diversity and leverage data from multiple languages. Furthermore, our multimodal time series forecasting library is built upon a simple integration framework that only uses a projection layer for fine-tuning LLMs. How to fine-tune LLMs more efficiently and effectively for time series analysis remains an interesting topic. Despite these limitations, we hope that the datasets and library we have constructed will facilitate broader research and applications in multimodal time series analysis.
Appendix C Statistics of Textual Data
@@
Appendix D Prompt designed for LLM preprocessing
@@
Appendix E Showcase of Raw and Preprocessed Text
We show the pre and post-processing text content of a report in Climate domain in Figure 9. We see LLM accurately summarized the factual content from original report.
Appendix F Manual Verification of LLM Preprocessing
@@
In the 8 fabricated instances, 5 were labeled with the data source as LLM, indicating that they can be filtered out. his results in a true hallucination rate of 3/127.
@@
Appendix G Details of MM-TSFlib
In terms of implementation, MM-TSFlib chooses to extend the widely used Time-Series Library (TSlib)444https://github.com/thuml/Time-Series-Library, thus ensuring ease of use. For LLM invocation, MM-TSFlib utilizes the popular and active Hugging Face555https://huggingface.co/models. For the projection layer, we use a multilayer perceptron (MLP) to keep it simple. To use Time-MMD dataset for TSF task, MM-TSFlib constrains the latest end date of the input text sequence to be earlier than the latest end date of the input sequence, in order to avoid information leakage. Overall, MM-TSFlib supports multimodal extensions of over 20 unimodal TSF algorithms via 7 open-source LLM models.
G.1 List of Supported TSF
@@
G.2 List of Supported LLM
For LLM, MM-TSFlib supports BERT[11], GPT-2[32] (Small, Medium, Large, Extra-Large), Llama-2-7B[37], and Llama-3-8B666https://llama.meta.com/llama3.
Appendix H More Details of Experimental Setup
H.1 Time-Series Forecasting Backbones
@@
DLibear, a linear model that performs forecasting by a direct regression upon historical time series with a one-layer linear model.

### v4

@@ 5 Potential Future Works
Beyond its efficacy in enhancing time-series forecasting accuracy through multimodality extension (Section 4), Time-MMD holds significant potential in advancing time series analysis across a wide spectrum. In the following section, we discuss how Time-MMD might transform conventional approaches, facilitate novel methodologies, and may broadly impact the time series analysis domain in future works.
@@
Multimodal Time-Series Imputation. Missing values in time series data, caused by sensor failures, system instability, or privacy concerns, pose a significant challenge in analysis. Conventional time-series imputation (TSI) methods (8; 19; 3; 4) often overlook valuable information captured in textual formats alongside the numerical data. For instance, incident reports, weather conditions, and special events can provide crucial context for imputing missing data points in traffic time series, but current methods fail to effectively incorporate this information. By constructing missing values using existing toolkits (16), Time-MMD can directly serve as multimodal imputation datasets. Time-MMD enables the integration of textual contextual information with numerical time series data, opening new avenues for multimodal time series analysis and enhancing imputation accuracy.
@@
Multimodal Time-Series Anomaly Detection. Detecting anomalies in time series data is crucial for identifying unusual patterns that may indicate faults, fraud, or other significant events (10; 5; 75). However, conventional anomaly detection methods (47; 7; 69; 53) are limited to expected pattern deviations in numerical data, overlooking valuable information in textual formats. For instance, news articles, social media posts, and market reports can provide critical context that influences financial market behavior and helps identify anomalies not evident from numerical data alone. Time-MMD enables a feasible way to support multimodal time series anomaly detection. For example, in influenza (health) dataset, if "flu outbreak" is mentioned in the text data, we can initially label the aligned numerical data as anomalous.
@@
Multimodal Foundation Time-Series Models. The introduction of Time-MMD, a comprehensive text-numeric TS dataset, is expected to significantly advance multimodal-based TS methods, including the development of multimodal foundation TS models beyond existing unimodal models (30). Time-MMD will facilitate further exploration with more complex and informative prompts, enhancing the performance and capabilities of fine-tuning methods. Additionally, it might spur research and development of multimodal models specifically tailored for TSA, an area with limited exploration compared to other domains like vision and video generation (54; 62; 68; 17).
@@ 7 Conclusion
Acknowledgements: This paper was supported in part by the NSF (Expeditions CCF-1918770, CAREER IIS-2028586, CAREER IIS-2144338, Medium IIS-1955883, Medium IIS-2106961, Medium IIS-2403240, PIPP CCF-2200269), CDC MInD program, Meta faculty gifts, and funds/computing resources from Georgia Tech and GTRI. This work used the Delta GPU Supercomputer at NCSA of UIUC through allocation CIS240288 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program.
@@ Appendix A Additional Related Dataset Work
Existing datasets [2, 71, 51, 18, 55, 15, 6] primarily focus on stock analysis tasks in the finance domain. Other multi-modal datasets for traffic demand prediction [51], news impact prediction [24] and electrocardiogram classification [34] are proposed recently. However, these datasets still do not address the aforementioned gaps. Especially constructing datasets for diverse domains, such as agriculture and security, is more challenging but holds substantial real-world impact.
Besides, some datasets focus on combining images with numerical sequences, where the images are typically sourced from spatial domains, such as satellite imagery. These works [21, 37] are centered on spatiotemporal data analysis, whereas our work focuses on time series analysis.
@@ Appendix B Limitations
Our work provides a comprehensive, high-quality multimodal time series dataset, but it still has limitations in terms of dataset diversity, as all the text data comes from English. We plan to extend Time-MMD to multilingual versions to better address diversity and leverage data from multiple languages [26]. In addition, utilizing Time-MMD to support other multimodal time series analysis tasks, such as multivariate time series forecasting and time series classification, requires additional dataset curation work. Moreover, considering multimodal time series datasets with images and audio is a worthwhile direction for future work. In this work, we chose to include text for the following reasons: (1) Text and numerical series commonly coexist across multiple domains. (2) With the rise of LLMs, an increasing number of time-series methods are incorporating text based on large language models. Collecting data that further align images [21] ,audio [31] or graph [39] with text and numeric is meaningful but beyond the scope of this work. Lastly, this work focuses on time series forecasting and does not cover interacting dynamical systems or equation discovery [46], where language could provide insights into the qualitative behavior of the system.
Our multimodal time series forecasting library is built upon a simple integration framework that only uses a projection layer for fine-tuning LLMs. How to fine-tune LLMs more efficiently and effectively for time series analysis remains an interesting topic. Moreover, addressing the issue of distributional shift [77, 79, 38] in the text should also be considered in future work to build robust models. In summary, our approach to achieve multimodal time series forecasting based on Time-MMD dataset is a first-cut reasonable approach rather than the optimal solution. There could be better ways to utilize it, such as Retrieval-Augmented Generation [20].
Appendix C Considerations for Domain Selection
Our domain selection is based on the following three considerations:
The remaining three domains, including social good, agriculture, and security, are often overlooked by researchers. We included these domains to contribute to a more comprehensive dataset.
Additionally, we also considered underrepresented groups, as reflected in the Health (Africa) dataset, compared with the Health (U.S.).
These domains exhibit differences in numerical data, textual data, and performances. To the best of our knowledge, our Time-MMD is the most domain-inclusive multimodal time-series dataset.
Appendix D Details of Data Sources and Variables by Domain
D.1 Agricultural Domain Dataset Sources
D.1.1 Numeric Data
The sequence raw data is sourced from the retail broiler composite, provided by the United States Department of Agriculture (USDA) Economic Research Service (ERS). The raw data can be accessed at the USDA ERS website555https://www.ers.usda.gov/data-products/meat-price-spreads/documentation/.
D.1.2 Text Data
Report: Source 1
Frequency: Weekly
Name: USDA Broiler Market News Report.
Description: This report provides a summary of the national broiler market, with data consolidated from a number of linked reports. The release contains price information for broiler parts by U.S. region as well as in Mexico City, slaughter estimates, availability, and national fowl market and miscellaneous poultry information.
Report: Source 2
Frequency: Daily
Name: Daily National Broiler Market at a Glance.
Description: This report summarizes the daily trends in the national broiler market, with commentary on price, demand and supply, market activity, and trade.
Search
Frequency: 10 results per week.
Keywords: United States Broiler Market; Chicken prices
D.2 Climate Domain Dataset Sources
The data is collected by the NOAA National Centers for Environmental Information and spans from 1895 to the present. It offers a comprehensive historical dataset for analyzing long-term climate trends and their implications for energy and environmental policies. The raw data can be accessed at the NOAA.
D.2.1 Numeric Data
Source
D.2.2 Text Data
Report: Source 1
Frequency: Month
Name: Drought Report of NOAA National Centers for Environmental Information666https://www.ncei.noaa.gov/access/monitoring/monthly-report/drought/200602.
Description: Drought Highlights;National Overview;Regional Overview
Report: Source 2
Frequency: Month
Name: National Climate Report of NOAA National Centers for Environmental Information777https://www.ncei.noaa.gov/access/monitoring/monthly-report/national/200602.
Description: National Overview;Monthly and Seasonal Highlights;
Search
Frequency: 10 results per week.
Keywords: Drought and extreme weather; Precipitation
D.3 Economy Domain Dataset Sources
D.3.1 Numeric Data
Source
D.3.2 Text Data
Report: Source 1
Frequency: Monthly
Name: U.S. International Trade in Goods and Services 888https://www.census.gov/foreign-trade/Press-Release/ft900_index.html.
Report: Source 2
Frequency: Monthly
Name: Advance Economic Indicators Report999https://www.census.gov/econ/indicators/current/index.html.
Search
Frequency: 10 results per week.
Keywords: International Trade Balance
D.4 Energy Domain Dataset Sources
This subsection outlines the sources and characteristics of the sequence and raw textual data that form the energy component of the series-text dataset. This domain specifically focuses on gasoline prices, which are crucial for analyzing economic stability, consumer spending patterns, and the overall energy market dynamics.
D.4.1 Numeric Data
The sequence raw data comprises weekly statistics of U.S. gasoline prices, measured in dollars per gallon. Monitoring these prices provides insights into energy market fluctuations and helps predict changes in economic conditions and policy adjustments.
Source and Accessibility
This data is collected by the U.S. Energy Information Administration (EIA) and spans from April 5, 1993, to the present, offering a comprehensive historical view of fuel economics. The raw data can be accessed at the EIA website101010https://www.eia.gov/petroleum/gasdiesel/.
D.4.2 Text Data
We collect raw text from various EIA publications and reports, as they provide contextual information and expert analysis related to the Numeric Data.
Report Source 1: Annual Energy Outlook
Frequency: Annually
Source: U.S. Energy Information Administration111111https://www.eia.gov/outlooks/aeo/archive.php.
Type: Annual Energy Outlook reports
Coverage: From 1979 to present
Report Source 2: Weekly Petroleum Status Report
Frequency: Weekly
Source: U.S. Energy Information Administration121212https://www.eia.gov/petroleum/supply/weekly/archive/.
Type: Weekly Petroleum Status Reports
Coverage: From 2011 to present
Search
Frequency: 10 results per week.
Keywords: Gasoline prices
D.5 Social Good Domain Dataset Sources
This subsection describes the sources and characteristics of the sequence and raw textual data that comprise the social good component of the series-text dataset. This domain focuses on unemployment statistics in the United States, segmented by various racial groups, reflecting the societal impact of economic disparities.
D.5.1 Numeric Data
The sequence raw data consists of monthly unemployment statistics for the United States, disaggregated by race. These statistics include data for the following racial groups: White, Black or African American, Asian, American Indian or Alaska Native, and Native Hawaiian or Other Pacific Islander.
This data is sourced from the United States Bureau of Labor Statistics (BLS) and spans from 1954 to the present, providing a long-term view of employment trends across different racial demographics. The raw data can be accessed at the BLS website131313https://data.bls.gov/home.htm.
D.5.2 Text Data
The raw textual data is sourced from official reports which provide a comprehensive analysis of employment conditions, enriched with context that is crucial for understanding the nuances of unemployment across different racial groups.
Report Source 1
Frequency: Monthly
Source: U.S. Department of Labor, BLS141414https://www.bls.gov/bls/news-release/empsit.htm.
Type: Official report on monthly employment situations
Coverage: From 1994 to present
Quantity: TBD
Relevance: Directly relevant (Siamese)
Report Source 2
Frequency: Annually
Source: U.S. Department of Labor, BLS151515https://www.bls.gov/opub/reports/race-and-ethnicity/2015/home.htm.
Type: Official report on annual labor force characteristics by race and ethnicity
Coverage: From 2015 to present
Quantity: TBD
Relevance: Directly relevant
Search
Frequency: 10 results per week.
Keywords: Unemployment rate; Employment situation
D.6 Public Health (United States) Domain Dataset Sources
This subsection outlines the sources and characteristics of the sequence and raw textual data that form the public health component of the series-text dataset. This domain specifically focuses on Influenza-Like Illness (ILI) statistics, which are crucial for monitoring seasonal and pandemic influenza outbreaks, guiding public health interventions, and planning healthcare resources effectively.
D.6.1 Numeric Data
The sequence raw data includes weekly statistics of ILI cases in the United States, sourced from the Centers for Disease Control and Prevention (CDC). ILI, representing a category of health conditions characterized by symptoms similar to those of influenza, serves as an important public health indicator for tracking influenza activity.
Source and Accessibility
The data is collected and made available by the CDC’s ILINet system, which has been tracking ILI cases since 1954. This extensive dataset allows for robust historical trend analysis and modeling of influenza patterns. The raw influenza patients data can be accessed at the CDC website161616https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html.
D.6.2 Text Data
Raw textual data in the public health domain includes official reports and related news, providing insights into the context and implications of the ILI statistics.
Report Source 1: Weekly U.S. Influenza Surveillance Report
Frequency: Weekly
Source: CDC171717https://www.cdc.gov/flu/weekly/pastreports.htm.
Type: Official Weekly U.S. Influenza Surveillance Report
Coverage: From 1999 to present
Report Source 2: Annual Flu Season Key Studies and News Reports
Frequency: Annually
Source: CDC181818https://www.cdc.gov/flu/spotlights/index.htm.
Type: Key studies and news reports from the flu season
Coverage: From 2019 to present
Search
Frequency: 10 results per week.
Keywords: influenza; epidemic
D.7 Public Health (Africa) Domain Dataset Sources
D.7.1 Numeric Data
Source:
D.7.2 Text Data
Report: Source 1
Frequency: Weekly
Name: Influenza Virological Surveillance in the WHO African Region202020https://www.afro.who.int/publications/influenza-virological-surveillance-who-african-region-epidemiological-week-39-2017.
Report: Source 2
Frequency: Daily
Name: Press of Africa Centres for Diseases Control212121https://africacdc.org/news-item/africa-centres-for-diseases-control-and-prevention-launches-new-networks-to-fight-health-threats-in-africa/.
Description:
Search
Frequency: 10 results per week.
Keywords: influenza; epidemic
D.8 Environment Domain Dataset Sources
D.8.1 Numeric Data
Source
D.8.2 Text Data
Report: Source 1
Frequency: Daily
Name: Press about Air Quality of Department of Environmental Conservation of New York222222https://dec.ny.gov/news/press-releases/2021/5/dec-directs-pvs-chemical-solutions-inc-to-temporarily-cease-operations.
Report: Source 2
Frequency: Daily
Name: Article about air quality from New York National Broadcasting Company232323https://www.nbcnewyork.com/tag/air-quality/.
Search
Frequency: 10 results per week.
Keywords: New York air quality; New York air pollution
D.9 Traffic Domain Dataset Sources
D.9.1 Numeric Data
Source
D.9.2 Text Data
Report: Source 1
Frequency: Weekly
Name: Weekly Traffic Volume Report242424https://datahub.transportation.gov/stories/s/Weekly-Traffic-Volume-Report/3g63-ik4u/.
The Traffic Volume report estimates the vehicle miles traveled (VMT) for interstate highways and how the total travel measured by VMT compares with travel that occurred in the same week of the previous year. The VMT is further split into passenger vehicle and truck components. The information gives new insights into the effect on traffic by storm activity, economic fluctuations, and other variables that could not be evaluated using the monthly report.
Search
Frequency: 10 results per week.
Keywords: Travel; Mobility
D.10 Security Domain Dataset Sources
D.10.1 Numeric Data
Source
D.10.2 Text Data
Report: Source 1
Frequency: Uncertain
Name: Billion-Dollar Weather and Climate Disasters252525https://www.ncei.noaa.gov/access/billions/.
Report: Source 2
Frequency: Uncertain
Name: Disaster and emergency declarations262626https://www.fema.gov/about/openfema/data-sets.
Appendix E Statistics of Textual Data
@@
Appendix F Prompt designed for LLM preprocessing
@@
Appendix G Showcase of Raw and Preprocessed Text
We show the pre and post-processing text content of a report in Climate domain in Figure 9. We observe that LLMs are able to accurately summarized the factual content from original report.
Appendix H Manual Verification of LLM Preprocessing
@@
In the 8 fabricated instances, 5 were labeled with the data source as LLM, indicating that they can be filtered out. This results in a true hallucination rate of 3/127.
@@
Appendix I Case Study of Semantic Alignment
A case example of textual data explaining numerical data changes is shown in Figure 10 from the Health dataset. As the reports and search data are collected while ensuring strong time alignments, their textual content can provide clear illustrations on numerical series changes. For example, during Feb 2008, the health care numerical data experienced a significant upward trend, while the summarized report data states "…… The proportion of deaths attributed to pneumonia and influenza was above the epidemic threshold for the fourth consecutive week ……". Explicit alignments like this can be easily found in our data.
Appendix J Details of MM-TSFlib
In terms of implementation, MM-TSFlib chooses to extend the widely used Time-Series Library (TSlib)272727https://github.com/thuml/Time-Series-Library, thus ensuring ease of use. For LLM invocation, MM-TSFlib utilizes the popular and active Hugging Face282828https://huggingface.co/models. For the projection layer, we use a multilayer perceptron (MLP) to keep it simple, apply instance normalization followed by adjusting the mean to the corresponding historical average to keep it rational. We tune the initialization values of the linear weighting mechanism to prevent overfitting or reduce learning difficulty. To use Time-MMD dataset for TSF task, MM-TSFlib constrains the latest end date of the input text sequence to be earlier than the latest end date of the input sequence, in order to avoid information leakage. Overall, MM-TSFlib supports multimodal extensions of over 20 unimodal TSF algorithms via 7 open-source LLM models.
J.1 List of Supported TSF
@@ Appendix I Case Study of Semantic Alignment
J.2 List of Supported LLM
For LLM, MM-TSFlib supports BERT [14], GPT-2 [50] (Small, Medium, Large, Extra-Large), Llama-2-7B [58], and Llama-3-8B292929https://llama.meta.com/llama3.
Appendix K More Details of Experimental Setup
K.1 Time-Series Forecasting Backbones
@@
DLinear, a linear model that performs forecasting by a direct regression upon historical time series with a one-layer linear model.